OpenStudy (anonymous):

How can you solve the equation y minus two point nine equals five point three? Subtract two point nine from both sides. Add five point three to both sides. Subtract five point three from both sides. Add two point nine to both sides.

OpenStudy (yttrium):

ooops! What do you think @iceicebaby ?

OpenStudy (anonymous):

You would add 5.3 to both sides.

OpenStudy (anonymous):

wow thanks for telling me i was trying to figure this out

OpenStudy (yttrium):

It's actually no right @iceicebaby , it must be the last choice.

OpenStudy (anonymous):

Do you understand how to do it though?

OpenStudy (anonymous):

add Add two point nine to both sides is what i was thinking but i was not sure

OpenStudy (yttrium):

well, you are actually correct @iceicebaby

OpenStudy (anonymous):

This would be your equation: y-2.9=5.3 You would need to flip the equation y=2.9+5.3 Then add 5.3 to both sides

OpenStudy (yttrium):

Huh? You actually did adding 2.9 to both sides. That's why it became y = 2.9 + 5.3 :) @GeckoBro